### v1

# Absence of shift-invariant Gibbs states (delocalisation) for one-dimensional

Abstract.
We show that a modification of the proof of our paper [2], in the spirit of [3], shows delocalisation in the long-range Discrete Gaussian Chain, and generalisations thereof, for any decay power $\alpha>2$ and at all temperatures. The argument proceeds by contradiction: any shift-invariant and localised measure (in the $L^{1}$ sense), is a convex combination of ergodic localised measures. But the latter cannot exist: on one hand, by the ergodic theorem, the average of the field over growing boxes would be almost surely bounded ; on the other hand the measure would be absolutely continuous with respect to its height-shifted translates, as a simple relative entropy computation shows. This leads to a contradiction and answers, in a non-quantitative way, an open question stated in a recent paper [7] of C. Garban.
Acknowledgments. We thank Christophe Garban for a helpful feedback.
Research of A.L.N. and L.C. have been supported by the CNRS IRP (International Research Project) EURANDOM “Random Graph, Statistical Mechanics and Networks” and by the LabEx PERSYVAL- Lab (ANR-11-LABX-0025-01) funded by the French program Investissement d’avenir. W.M.R. is funded by Vidi grant VI.Vidi.213.112 from the Dutch Research Council.
1. Introduction
The long-range Discrete Gaussian Chain (DGC) is a $\mathbb{Z}$-valued version of the $1d$ long-range Ising model, where the height variables $\phi_{i}\in\mathbb{Z}$ at site $i\in\mathbb{Z}$ form a height configuration $\phi$ in the state-space $\mathbb{Z}^{\mathbb{Z}}$, which is sampled according to the Gibbs measure with formal Hamiltonian
$H(\phi):=\sum_{i\neq j\in\mathbb{Z}}J_{\alpha}(|i-j|)\cdot|\phi_{i}-\phi_{j}|^{2}$ (1)
with long-range coupling constants $J_{\alpha}(k)\sim k^{-\alpha}$, the most classical choice being $J_{\alpha}(k)=k^{-\alpha}$. It has been introduced as an effective model for an interface between two ordered phases in $2d$ long-range Ising models with the same polynomially decaying coupling constants, and was studied in [9, 6, 12, 13, 2].
The localisation/delocalisation of models with Hamiltonians (1) is analogous to the same question for $2d$ nearest neighbour integer-valued GFF (also called Discrete Gaussian model) with $\phi:\mathbb{Z}^{2}\longrightarrow\mathbb{Z}$ whose delocalisation at high temperature has been proved in [5, 10], see also [11].
The behaviour of the long-range DGC with Hamiltonian (1) is as follows:
  • For $\alpha\in(1,2)$ the field $\phi$ is localised at any inverse temperature $\beta$, as recently proved in [7].
  • For $\alpha=2$, there is a localisation/delocalisation phase transition, which was originally proved by Fröhlich and Zegarlinski [6]. More precisely, the localisation at low temperature was proved in [6] and the delocalisation at high temperature was proved in [9, 6].
  • For $\alpha>2$, Fröhlich and Zegarlinski [6] conjecture that fluctuations are of order $n^{\min\{\alpha-2,1\}/2}$ at all temperatures, and give a heuristical energy-entropy argument.
  • For $\alpha\in[2,3)$, the paper [7] proposes a new proof of delocalisation at high temperature, relying on [5], which goes much beyond the above conjecture and exhibits the phenomenon of ”invisibility of the integers”: an invariance principle towards a $H$-fractional Brownian motion of Hurst index $\frac{\alpha-2}{2}$ holds, and this limiting Gaussian field is the same as the one arising from the real-valued Gaussian Chain. The diffusivity constant is also computed.
  • For $\alpha>3$, there is delocalisation at all temperatures and fluctuations follow a standard central limit scaling of order $n^{1/2}$, see [7].
The following open problem is stated in [7] as Open Problem 5:
Prove the delocalisation of the discrete Gaussian chain at all temperatures when $\alpha\in(2,3)$, possibly with quantitative estimates on the variance. A heuristics is discussed in [6], but as far as we know, a proof is still missing.
In the work [2], the four present authors studied the $2d$ long-range Ising model, and proved absence of Dobrushin states (extremal non-transation invariant Gibbs measures arising from Dobrushin boundary conditions) for any $\alpha>2$ and at any temperatures, by means of Araki’s relative entropy method from [1, 3, 4].
The goal of the present note is to show that the same method leads to a (non-quantitative) proof of delocalisation of the long-range DGC or SOS model, and generalisations thereof, at all temperatures for any $\alpha>2$. Quantitative estimates on the delocalisation should be derivable, providing some bounds on the fluctuation exponent, which is conjectured in [6] to be equal to $\min\{\alpha-2,1\}/2$ as stated above, but in this short note we restrict ourselves to the following rigorous qualitative statement.
2. Non-quantitative proof of delocalisation
Consider $\mathbb{Z}$-valued fields on $\mathbb{Z}$ under potential $V(x)=|x|^{p}$ with $p\in[1,2]$ and long-range polynomially decaying interactions $J_{\alpha}(x)\sim|x|^{-\alpha}$. The Hamiltonian corresponding to this class of models is :
$H(\phi)=-\sum_{i\neq j\in\mathbb{Z}}J_{ij}V(\phi_{i}-\phi_{j}),$ (2)
where $J_{ij}=J_{\alpha}(|i-j|)$ (positivity is not assumed) and $\phi_{i}$ takes values in $\mathbb{Z}$. The model makes sense for $\alpha>1$. The DGC corresponds to $V(x)=x^{2}$, and the SOS model corresponds to $V(x)=|x|$. For general $p$, the model is also called $p$-SOS.
Let $\mu^{\omega}_{\Lambda}$ denote the finite-volume measure with Hamiltonian (2) with height-$\omega$ boundary condition outside the box $\Lambda\Subset\mathbb{Z}$. More precisely, for $\beta>0$,
$\mu^{\omega}_{\Lambda}(\phi)\propto\exp\left(-\beta H(\phi)\right)1_{\phi_{i}\in\mathbb{Z}\;\forall i\text{ and }\phi_{i}=\omega\;\forall i\in\Lambda^{c}}.$
In the case $\omega\equiv 0$ we write $\mu^{0}_{\Lambda}$. A Gibbs measure for the Hamiltonian (2) is a probability measure $\mu$ on $\Omega=\mathbb{Z}^{\mathbb{Z}}$ such that for any $\Lambda\Subset\mathbb{Z}$, we have
[…]
Expectations under a measure $\nu$ are written $\langle\cdot\rangle_{\nu}$. We say that a measure $\nu$ is shift-invariant if $\nu(\phi)=\nu(\tau(\phi))$ for any translation $\tau:z\mapsto z+v$, $v\in\mathbb{Z}$. The set of shift-invariant Gibbs measures is denoted by $\mathcal{G}_{si}$.
We write $\nu_{t}(\phi)=\nu(\phi+t)$, when we shift the heights by an amount $t$.
Theorem 1.
For any $\beta>0$ and $\alpha>2$, delocalisation holds for the models with Hamiltonian (2) in the following sense:
  1. (1)
    Let $p\in\{1,2\}$, i.e. consider either the SOS model or the DGC.
    There does not exist any shift-invariant Gibbs measure $\mu\in\mathcal{G}_{si}$ such that
    ${\langle|\phi_{0}|\rangle}_{\mu}<C\text{ for some finite constant }C.$ (3)
  2. (2)
    Let $p\in(1,2)$. There does not exist any shift-invariant Gibbs measure $\mu\in\mathcal{G}_{si}$ such that
    ${\langle\phi_{0}^{2}\rangle}_{\mu}<C\text{ for some finite constant }C.$ (4)
Corollary 1.
For any $\beta>0$ and $\alpha>2$, there is no sequence of boxes $(\Lambda_{n})_{n}\uparrow\mathbb{Z}$ such that
$\langle\phi_{0}^{2}\rangle^{0}_{\Lambda_{n}}<C\quad\text{ for every }n\in\mathbb{N},$
and such that the sequence ${(\mu^{0}_{\Lambda_{n}})}_{n}$ weakly converges to a shift-invariant111or even shift periodic or shift quasiperiodic Gibbs measure.
Corollary 1 easily follows from Fatou’s Lemma and Theorem 1. The result should be true without the assumption of shift-invariance of the limit, but we cannot exclude any (subsequential) weak limit of the $\mu^{0}_{\Lambda_{n}}$ not to be extremal, nor exclude it to be a mixture of exclusively non-shift invariant extremal measures. Such exclusions can be proved in the case of planar interactions, see the thesis of Sheffield [12].
Note that Lammers and Ott proved the FKG inequality for the absolute value of the field (for super-Gaussian potentials, i.e. $p\in[1,2]$ in our model), which implies, in the context of planar interactions, the equivalence of the two notions of delocalisation, namely delocalisation in finite volume and delocalisation of shift-invariant Gibbs measures. See [11] Theorem 2.7.
Proof of Theorem 1.
The argument is based on [3, 2], and proceeds by contradiction: if we suppose $\mu$ is shift-invariant and localised ((3) holds), then it is a convex combination of ergodic localised measures. But an ergodic localised measure $\nu$ cannot exist: on one hand, the average of the field over growing boxes would be $\nu$ almost surely bounded by ergodicity and by our under $L^{1}$ assumption (3); on the other hand $\nu$ would be absolutely continuous with respect to its vertical translate $\nu_{t}$ as a simple relative entropy computation shows. This leads to a contradiction for $t$ large enough.
More precisely, fix $\beta>0$ and suppose there exists $\mu\in\mathcal{G}_{si}$ which is localised:
$\exists C\in[0,\infty)\text{ such that }\langle|\phi_{0}|\rangle_{\mu}<C.$ (5)
Then $\mu$ has a unique decomposition $w_{\mu}$ onto shift-invariant extremal Gibbs measures (see [8], Theorem 14.17):
$\mu=\int_{ex\mathcal{G}_{si}}\nu\;w_{\mu}(d\nu).$
Moreover, this decomposition is supported on localised measures. Indeed, as $\langle|\phi_{0}|\rangle_{\mu}=\int_{ex\mathcal{G}_{si}}{\langle|\phi_{0}|\rangle}_{\nu}w_{\mu}(d\nu)<C,$ then for $w_{\mu}$ almost every $\nu\in ex\mathcal{G}_{si}$, there exists some $C_{\nu}<\infty$ such that
${\langle|\phi_{0}|\rangle}_{\nu}\leq C_{\nu}.$ (6)
We now prove that an extremal shift-invariant localised Gibbs measure $\nu$ for the Hamiltonian (2) cannot exist if $\alpha>2$.
Assume $\nu\in ex\mathcal{G}_{si}$ such that (6) holds.
On one hand, any $\nu\in ex\mathcal{G}_{si}$ is ergodic (see [8] Theorem 14.15) and thus the field is strongly localised under $\nu$, in the following sense.
Let $M_{n}=\frac{1}{2n}\sum_{|i|\leq n}\phi_{i}$. By the ergodic theorem, under hypothesis (3),
$\nu\big{(}M_{n}\to{\langle\phi_{0}\rangle}_{\nu}\big{)}=1,$ (7)
and thus for any $t\in\mathbb{Z}$,
$\nu_{t}\big{(}M_{n}\to{\langle\phi_{0}\rangle}_{\nu}+t\big{)}=1.$ (8)
On the other hand, the relative entropy between $\nu$ and $\nu_{t}$ is bounded if $\alpha>2$. To prove this, we follow Section 4 of [3] with a simplified local transformation of the measure $\nu$. Recall that the relative entropy between two measures $\mu$ and $\nu$, denoted $RE(\mu|\nu)$, is defined to be
$RE(\mu|\nu):={\langle\log({d\mu}/{d\nu})\rangle}_{\mu}$
when $\mu\ll\nu$, and +$\infty$+ ∞ otherwise. Let $t\in\mathbb{N}$ be fixed, and define ”the step of height $t$” as
$a_{i}^{n}(t)$ =$=\left\{\begin{array}[]{ll}t&\text{if }|i|<n\\ 0&\text{if }|i|\geq n.\end{array}\right.$
Define the transformation $T_{t,n}$ (”adding the step of height $t$ in a box of size $n$”) on $\Omega=\mathbb{Z}^{\mathbb{Z}}$ by setting
$(T_{t,n}\phi)_{i}=\phi_{i}+a_{i}^{n}(t).$
Let $\nu_{t,n}$ be the transformed Gibbs measure defined by
${\langle f\rangle}_{\nu_{t,n}}=\int\nu(d\phi)f(T_{t,n}^{-1}\phi).$
Then $\nu_{t,n}$ is absolutely continuous with respect to $\nu$, with Radon-Nykodim derivative given by222this follows from the fact that $\nu$ is a Gibbs measure.
$\frac{d\nu_{t,n}}{d\nu}(\phi)=\exp(-\beta\sum_{i\neq j\in\mathbb{Z}}J_{ij}(V(T_{t,n}\phi_{i}-T_{t,n}\phi_{j})-V(\phi_{i}-\phi_{j}))).$
To compute the relative entropy, note that
${\langle V(T_{t,n}\phi_{i}-T_{t,n}\phi_{j})-V(\phi_{i}-\phi_{j})\rangle}_{\nu}$
$={\langle|\phi_{i}+a_{i}^{n}(t)-\phi_{j}-a_{j}^{n}(t)|^{p}-|\phi_{i}-\phi_{j}|^{p}\rangle}_{\nu}.$ (9)
In the case $p=2$, we get
($=2(a_{i}^{n}(t)-a_{j}^{n}(t)){\langle\phi_{i}-\phi_{j}\rangle}_{\nu}+(a_{i}^{n}(t)-a_{j}^{n}(t))^{2}=(a_{i}^{n}(t)-a_{j}^{n}(t))^{2}.$
In the cases $p\in[1,2)$, we have the upper bound333using $|a+b|^{p}\leq 2^{p-1}(|a|^{p}+|b|^{p})$ for any $p\geq 1$.
(2)$\leq(2^{p-1}-1){\langle|\phi_{i}-\phi_{j}|^{p}\rangle}_{\nu}+2^{p-1}|a_{i}^{n}(t)-a_{j}^{n}(t)|^{p}.$
The first term vanishes in the case $p=1$, and is bounded in the remaining cases $p\in(1,2)$ under hypothesis (4), as $L^{p}(\nu)\subset L^{2}(\nu)$.
Altogether, assuming (3) for $p\in\{1,2\}$ and (4) for $p\in(1,2)$, we have
$RE(\nu|\nu_{t,n})$ $=\big{\langle}\sum_{i\neq j\in\mathbb{Z}}J_{ij}(V(T_{t,n}\phi_{i}-T_{t,n}\phi_{j})-V(\phi_{i}-\phi_{j})){\big{\rangle}}_{\nu}$
$\leq\sum_{|i|\leq n}\sum_{j\in\mathbb{Z}}|J_{ij}|(C_{1}+C_{2}(a_{i}^{n}(t)-a_{j}^{n}(t))^{p})$
$=\sum_{|i|\leq n}\sum_{|j|>n}|J_{ij}|(C_{1}+C_{2}t^{p})\leq(C_{3}+C_{4}n^{2-\alpha})(C_{1}+C_{2}t^{p})$
$<C_{5}+C_{6}t^{p}\text{ if }\alpha>2.$
where the above constants are positive and finite, they may depend on $p$ but do not depend on $n$.
[…]
$RE(\nu|\nu_{t})\leq\lim_{n\to\infty}RE(\nu|\nu_{t,n})$ $<C_{5}+C_{6}t^{p}\text{ if }\alpha>2.$
We deduce that, for any $t\in\mathbb{Z}$, $\nu$ and $\nu_{t}$ are absolutely continuous with respect to each other if $\alpha>2$.
This is in contradiction with the strong localisation property of ergodic measures ((7) and (8) imply that for any $t\neq 0$ the measures $\nu$ and $\nu_{t}$ are singular with respect to each other).