### v2

# Absence of shift-invariant Gibbs states (delocalisation) for one-dimensional

Abstract.
We show that a modification of the proof of our paper [2], in the spirit of [3], shows delocalisation in the long-range Discrete Gaussian Chain, and generalisations thereof, for any decay power $\alpha>2$ and at all temperatures. The argument proceeds by contradiction: any shift-invariant and localised measure (in the $L^{1}$ sense), is a convex combination of ergodic localised measures. But the latter cannot exist: on one hand, by the ergodic theorem, the average of the field over growing boxes would be almost surely bounded ; on the other hand the measure would be absolutely continuous with respect to its height-shifted translates, as a simple relative entropy computation shows. This leads to a contradiction and answers, in a non-quantitative way, an open question stated in a recent paper of C. Garban [7].
Acknowledgments. We thank Christophe Garban for a helpful feedback.
Research of A.L.N. and L.C. have been supported by the CNRS IRP (International Research Project) EURANDOM “Random Graph, Statistical Mechanics and Networks” and by the LabEx PERSYVAL- Lab (ANR-11-LABX-0025-01) funded by the French program Investissement d’avenir. W.M.R. is funded by Vidi grant VI.Vidi.213.112 from the Dutch Research Council.
1. Introduction
The long-range Discrete Gaussian Chain (DGC) is a $\mathbb{Z}$-valued version of the $1d$ long-range Ising model, where the height variables $\phi_{i}\in\mathbb{Z}$ at site $i\in\mathbb{Z}$ form a height configuration $\phi$ in the state-space $\mathbb{Z}^{\mathbb{Z}}$, which is sampled according to the Gibbs measure with formal Hamiltonian
$H(\phi):=\sum_{i\neq j\in\mathbb{Z}}J_{\alpha}(|i-j|)\cdot|\phi_{i}-\phi_{j}|^{2}$ (1)
with long-range coupling constants $J_{\alpha}(k)\sim k^{-\alpha}$, the most classical choice being $J_{\alpha}(k)=k^{-\alpha}$. It has been introduced as an effective model for an interface between two ordered phases in $2d$ long-range Ising models with the same polynomially decaying coupling constants, and was studied in [9, 6, 12, 13, 2].
The localisation/delocalisation of models with Hamiltonians (1) is analogous to the same question for $2d$ nearest neighbour integer-valued GFF (also called Discrete Gaussian model) with $\phi:\mathbb{Z}^{2}\longrightarrow\mathbb{Z}$ whose delocalisation at high temperature has been proved in [5, 10], see also [11].
The behaviour of the long-range DGC with Hamiltonian (1) is as follows:
  • For $\alpha\in(1,2)$ the field $\phi$ is localised at any inverse temperature $\beta$, as recently proved in [7].
  • For $\alpha=2$, there is a localisation/delocalisation phase transition, which was originally proved by Fröhlich and Zegarlinski [6]. More precisely, the localisation at low temperature was proved in [6] and the delocalisation at high temperature was proved in [9, 6].
  • For $\alpha>2$, Fröhlich and Zegarlinski [6] conjecture that fluctuations are of order $n^{\min\{\alpha-2,1\}/2}$ at all temperatures, and give a heuristical energy-entropy argument.
  • For $\alpha\in[2,3)$, the paper [7] proposes a new proof of delocalisation at high temperature, relying on [5], which goes much beyond the above conjecture and exhibits the phenomenon of ”invisibility of the integers”: an invariance principle towards a $H$-fractional Brownian motion of Hurst index $\frac{\alpha-2}{2}$ holds, and this limiting Gaussian field is the same as the one arising from the real-valued Gaussian Chain. The diffusivity constant is also computed.
  • For $\alpha>3$, there is delocalisation at all temperatures and fluctuations follow a standard central limit scaling of order $n^{1/2}$, see [7].
The following open problem is stated in [7] as Open Problem 5:
Prove the delocalisation of the discrete Gaussian chain at all temperatures when $\alpha\in(2,3)$, possibly with quantitative estimates on the variance. A heuristics is discussed in [6], but as far as we know, a proof is still missing.
In the work [2], the four present authors studied the $2d$ long-range Ising model, and proved absence of Dobrushin states (extremal non-transation invariant Gibbs measures arising from Dobrushin boundary conditions) for any $\alpha>2$ and at any temperatures, by means of Araki’s relative entropy method from [1, 3, 4].
The goal of the present note is to show that the same method leads to a (non-quantitative) proof of delocalisation of the long-range DGC or SOS model, and generalisations thereof, at all temperatures for any $\alpha>2$. Quantitative estimates on the delocalisation should be derivable, providing some bounds on the fluctuation exponent, which is conjectured in [6] to be equal to $\min\{\alpha-2,1\}/2$ as stated above, but in this short note we restrict ourselves to the following rigorous qualitative statement.
Note that such a simple proof would (and should) not work in the two-dimensional case, because the relative entropy between the field and its vertical translate in a box of size $n$ by $n$ will grow as fast as the perimeter of the box [^1] The constant $C_{3}$ in Equation (2) of the proof below would be replaced by $C_{3}n$., which is not bounded in dimension 2. This is consistent with the fact that at low temperature, the $2d$ nearest neighbor SOS model localises, so the long-range models should be localised also (even more by inequalities). Delocalisation at high temperature when $\alpha>4$ is the plausible conjecture (see [7], Open problem 7) but we expect the proof to be much harder.
2. Non-quantitative proof of delocalisation
Consider $\mathbb{Z}$-valued fields on $\mathbb{Z}$ under potential $V(x)=|x|^{p}$ with $p\in[1,2]$ and long-range polynomially decaying interactions $J_{\alpha}(x)\sim|x|^{-\alpha}$. The Hamiltonian corresponding to this class of models is :
$H(\phi)=-\sum_{i\neq j\in\mathbb{Z}}J_{ij}V(\phi_{i}-\phi_{j}),$ (2)
where $J_{ij}=J_{\alpha}(|i-j|)$ (positivity is not assumed) and $\phi_{i}$ takes values in $\mathbb{Z}$. The model makes sense for $\alpha>1$. The DGC corresponds to $V(x)=x^{2}$, and the SOS model corresponds to $V(x)=|x|$. For general $p$, the model is also called $p$-SOS.
Let $\mu^{\omega}_{\Lambda}$ denote the finite-volume measure with Hamiltonian (2) with height-$\omega$ boundary condition outside the box $\Lambda\Subset\mathbb{Z}$. More precisely, for $\beta>0$,
$\mu^{\omega}_{\Lambda}(\phi)\propto\exp\left(-\beta H(\phi)\right)1_{\phi_{i}\in\mathbb{Z}\;\forall i\text{ and }\phi_{i}=\omega\;\forall i\in\Lambda^{c}}.$
In the case $\omega\equiv 0$ we write $\mu^{0}_{\Lambda}$. A Gibbs measure for the Hamiltonian (2) is a probability measure $\mu$ on $\Omega=\mathbb{Z}^{\mathbb{Z}}$ such that for any $\Lambda\Subset\mathbb{Z}$, we have
[…]
Expectations under a measure $\nu$ are written $\langle\cdot\rangle_{\nu}$. We say that a measure $\nu$ is shift-invariant if $\nu(\phi)=\nu(\tau(\phi))$ for any translation $\tau:z\mapsto z+v$, $v\in\mathbb{Z}$. The set of shift-invariant Gibbs measures is denoted by $\mathcal{G}_{si}$.
We write $\nu_{t}(\phi)=\nu(\phi+t)$, when we shift the heights by an amount $t$.
Theorem 1.
For any $\beta>0$ and $\alpha>2$, delocalisation holds for the models with Hamiltonian (2) in the following sense:
  1. (1)
    Let $p\in\{1,2\}$, i.e. consider either the SOS model or the DGC.
    There does not exist any shift-invariant Gibbs measure $\mu\in\mathcal{G}_{si}$ such that localisation holds in the $L^{1}$ sense
    ${\langle|\phi_{0}|\rangle}_{\mu}<{\infty}.$ (4)
  2. (2)
    Let $p\in(1,2)$. There does not exist any shift-invariant Gibbs measure $\mu\in\mathcal{G}_{si}$ such that localisation holds in the $L^{2}$ sense
    ${\langle\phi_{0}^{2}\rangle}_{\mu}<{\infty}.$ (5)
Corollary 1.
For any $\beta>0$, $\alpha>2$, also finite-volume delocalisation holds for the ferromagnetic models (i.e. $J_{ij}\geq 0$) with Hamiltonian (2), in the following sense :
For any sequence of boxes $(\Lambda_{n})_{n}\uparrow\mathbb{Z}$ it holds
${\langle\phi_{0}^{2}\rangle}^{0}_{\Lambda_{n}}\to\infty\quad\text{ as }n\to\infty.$ (6)
Proof of Corollary 1.
Lammers and Ott proved the FKG inequality for the absolute value of the field (see [11], Theorem 2.8) which implies, in the context of planar interactions, the equivalence of the two notions of delocalisation, namely (D1) delocalisation in finite volume in the sense of (6), and (D2) absence of shift-invariant Gibbs measures (see [11], Theorem 2.7).
We notice that their proof of the implication (D2$\Rightarrow$(D1) holds without the planarity assumption on the interactions.
Indeed, following Section 7 of [11], one notices that $\phi_{i}=0\Leftrightarrow|\phi_{i}|=0$. The FKG lattice condition for the absolute value of the field (for which zero is minimal) and the DLR property (3) imply that for any $\Lambda\subset\Lambda^{\prime}\subseteq\mathbb{Z}$, the stochastic domination $|\phi|\preceq|\phi^{\prime}|$ holds under $\mu^{0}_{\Lambda}$ and $\mu^{0}_{\Lambda^{\prime}}$. In particular, the variances $\langle\phi_{i}^{2}\rangle^{0}_{\Lambda_{n}}=\langle|\phi_{i}|^{2}\rangle^{0}_{\Lambda_{n}}$ are increasing in $n$, for any $i\in\mathbb{Z}$, and the limiting variances
$\lim_{n\to\infty}{\langle\phi_{i}^{2}\rangle}^{0}_{\Lambda_{n}}$
exist in $[0,\infty]$ and coincide for all $i\in\mathbb{Z}$.
Assume the limiting variance is finite. We can thus follow [11] to construct a shift-invariant localised Gibbs measure, which cannot exist by Theorem 1, and leads to a contradiction, proving Corollary 1. ∎
Note that we do not exclude any (subsequential) weak limit of the $\mu^{0}_{\Lambda_{n}}$ not to be extremal, nor exclude it to be a mixture of exclusively non-shift invariant extremal measures. Such exclusions can be proved in the case of planar interactions, see the thesis of Sheffield [12].
Proof of Theorem 1.
The argument is based on [3, 2], and proceeds by contradiction: if we suppose $\mu$ is shift-invariant and localised ((4) holds), then it is a convex combination of ergodic localised measures. But an ergodic localised measure $\nu$ cannot exist: on one hand, the average of the field over growing boxes would be $\nu$ almost surely bounded by ergodicity and by our under $L^{1}$ assumption (4); on the other hand $\nu$ would be absolutely continuous with respect to its vertical translate $\nu_{t}$ as a simple relative entropy computation shows. This leads to a contradiction for $t$ large enough.
More precisely, fix $\beta>0$ and suppose there exists $\mu\in\mathcal{G}_{si}$ which is localised:
$\exists C\in[0,\infty)\text{ such that }\langle|\phi_{0}|\rangle_{\mu}<C.$ (7)
Then $\mu$ has a unique decomposition $w_{\mu}$ onto shift-invariant extremal Gibbs measures (see [8], Theorem 14.17):
$\mu=\int_{ex\mathcal{G}_{si}}\nu\;w_{\mu}(d\nu).$
Moreover, this decomposition is supported on localised measures. Indeed, as $\langle|\phi_{0}|\rangle_{\mu}=\int_{ex\mathcal{G}_{si}}{\langle|\phi_{0}|\rangle}_{\nu}w_{\mu}(d\nu)<C,$ then for $w_{\mu}$ almost every $\nu\in ex\mathcal{G}_{si}$, there exists some $C_{\nu}<\infty$ such that
${\langle|\phi_{0}|\rangle}_{\nu}\leq C_{\nu}.$ (8)
We now prove that an extremal shift-invariant localised Gibbs measure $\nu$ for the Hamiltonian (2) cannot exist if $\alpha>2$.
Assume $\nu\in ex\mathcal{G}_{si}$ is such that (8) holds.
On the one hand, any $\nu\in ex\mathcal{G}_{si}$ is ergodic (see [8] Theorem 14.15) and thus the field is strongly localised under $\nu$, in the following sense.
Let $M_{n}=\frac{1}{2n}\sum_{|i|\leq n}\phi_{i}$. By the ergodic theorem, under hypothesis (4),
$\nu\big{(}M_{n}\to{\langle\phi_{0}\rangle}_{\nu}\big{)}=1,$ (9)
and thus for any $t\in\mathbb{Z}$,
$\nu_{t}\big{(}M_{n}\to{\langle\phi_{0}\rangle}_{\nu}+t\big{)}=1.$ (10)
In particular, for any $t\neq 0$ the measures $\nu$ and $\nu_{t}$ are singular with respect to each other:
$\forall t\neq 0,\quad\nu\perp\nu_{t}.$ (11)
On the other hand, the relative entropy between $\nu$ and $\nu_{t}$ is bounded if $\alpha>2$. To prove this, we follow Section 4 of [3] with a simplified local transformation of the measure $\nu$. Recall that the relative entropy between two measures $\mu$ and $\nu$, denoted $RE(\mu|\nu)$, is defined to be
$RE(\mu|\nu):={\langle\log({d\mu}/{d\nu})\rangle}_{\mu}$
when $\mu\ll\nu$, and +$\infty$+ ∞ otherwise. Let $t\in\mathbb{N}$ be fixed, and define ”the step of height $t$” as
$a_{i}^{n}(t)$ =$=\left\{\begin{array}[]{ll}t&\text{if }|i|<n\\ 0&\text{if }|i|\geq n.\end{array}\right.$
Define the transformation $T_{t,n}$ (”adding the step of height $t$ in a box of size $n$”) on $\Omega=\mathbb{Z}^{\mathbb{Z}}$ by setting
$(T_{t,n}\phi)_{i}=\phi_{i}+a_{i}^{n}(t).$
Let $\nu_{t,n}$ be the transformed Gibbs measure defined by
${\langle f\rangle}_{\nu_{t,n}}=\int\nu(d\phi)f(T_{t,n}^{-1}\phi).$
Then $\nu_{t,n}$ is absolutely continuous with respect to $\nu$, with Radon-Nykodim derivative given by222this follows from the fact that $\nu$ is a Gibbs measure.
$\frac{d\nu_{t,n}}{d\nu}(\phi)=\exp(-\beta\sum_{i\neq j\in\mathbb{Z}}J_{ij}(V(T_{t,n}\phi_{i}-T_{t,n}\phi_{j})-V(\phi_{i}-\phi_{j}))).$
To compute the relative entropy, note that
${\langle V(T_{t,n}\phi_{i}-T_{t,n}\phi_{j})-V(\phi_{i}-\phi_{j})\rangle}_{\nu}$
$={\langle|\phi_{i}+a_{i}^{n}(t)-\phi_{j}-a_{j}^{n}(t)|^{p}-|\phi_{i}-\phi_{j}|^{p}\rangle}_{\nu}.$ (12)
In the case $p=2$, we get
($=2(a_{i}^{n}(t)-a_{j}^{n}(t)){\langle\phi_{i}-\phi_{j}\rangle}_{\nu}+(a_{i}^{n}(t)-a_{j}^{n}(t))^{2}=(a_{i}^{n}(t)-a_{j}^{n}(t))^{2}.$
In the cases $p\in[1,2)$, we have the upper bound333using $|a+b|^{p}\leq 2^{p-1}(|a|^{p}+|b|^{p})$ for any $p\geq 1$.
(2)$\leq(2^{p-1}-1){\langle|\phi_{i}-\phi_{j}|^{p}\rangle}_{\nu}+2^{p-1}|a_{i}^{n}(t)-a_{j}^{n}(t)|^{p}.$
The first term vanishes in the case $p=1$, and is bounded in the remaining cases $p\in(1,2)$ under hypothesis (5), as $L^{2}(\nu)\subset L^{p}(\nu)$.
Altogether, assuming (4) for $p\in\{1,2\}$ and (5) for $p\in(1,2)$, we have
$RE(\nu|\nu_{t,n})$ $=\big{\langle}\sum_{i\neq j\in\mathbb{Z}}J_{ij}(V(T_{t,n}\phi_{i}-T_{t,n}\phi_{j})-V(\phi_{i}-\phi_{j})){\big{\rangle}}_{\nu}$
$\leq\sum_{|i|\leq n}\sum_{j\in\mathbb{Z}}|J_{ij}|(C_{1}+C_{2}(a_{i}^{n}(t)-a_{j}^{n}(t))^{p})$
$=\sum_{|i|\leq n}\sum_{|j|>n}|J_{ij}|(C_{1}+C_{2}t^{p})\leq(C_{3}+C_{4}n^{2-\alpha})(C_{1}+C_{2}t^{p})$ (13)
$<C_{5}+C_{6}t^{p}\text{ if }\alpha>2.$
where the above constants are positive and finite, they may depend on $p$ but do not depend on $n$.
[…]
$RE(\nu|\nu_{t})\leq\lim_{n\to\infty}RE(\nu|\nu_{t,n})$ $<C_{5}+C_{6}t^{p}\text{ if }\alpha>2.$
We deduce that, for any $t\in\mathbb{Z}$, $\nu$ and $\nu_{t}$ are absolutely continuous with respect to each other if $\alpha>2$. This is in contradiction with (11) and proves the Theorem. ∎